OpenStudy (anonymous):

Simplify 6^6 ÷ 6^5 using the laws of exponents. Verify your answer by expanding the powers, then dividing. Show your work.

OpenStudy (saifoo.khan):

\[\Large a^b \div a^c = a^{b-c}\]

OpenStudy (saifoo.khan):

Using that law will give us something like: \[\Large 6^6 \div 6^5 = 6^{6-5} = 6^1\]

OpenStudy (anonymous):

is that 6^6-5 6 to the power of 6 - 5

OpenStudy (saifoo.khan):

No. 6-5 is in the exponent.

OpenStudy (anonymous):

thankyou